solve each system by substitution. y=-3x+19
y=2x-1

Answer by rfer(16322)

y=-3x+19
y=2x-1
2x-1=-3x+19
5x=20
x=4
---------------
y=2(4)-1
y=7
---------------
7=-3(4)+19
7=-12+19